Shift report 20020111 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Jan 11 2002 - 23:53:35 EST

  • Next message: Apache: "Shift report 20020112 00:00-08:00"

    ------------------------------------------------------------
    Date: Fri Jan 11 23:53:35 2002, Shift: 16:00-24:00
    Supervisor: IGB, with: 
    Log book#12 pages 47 - 48
    
    Shift Summary:
    ------------------------------------------------------------
    Machine Status:
    A few attempts a beam failed, but have had good beams from 2130.
    
    Achievements:
    
    Took 460K trigger 6 (FS) with FS at 4 degrees, 1/5(A) field, 
    even though I was only supposed to take 300K.  
    Change field settings in FS to 1/3 (A) and have nearly 100K trigger 6 after 1/2 hour.
    Took ca. 80K (out of 400) trigger 3 with MRS@45 degrees, 4A setting.
    Detector Problems:
    None detected
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 14 2002 - 11:31:42 EST